A pair of 7 sided dice are tossed. What is the probability that at least one of the dice has a value greater than or equal to 2?
Answer #1


When a pair of 7 sided dice are tossed, total possible outcomes are 7^2 = 49

Consider an event such that both the dice has value less than 2 i.e. 1 and there is only one such occurrance. In short, we do not need (1,1) as outcome.

Hence required probability is 1 - 1/49 = 48/49